Please Help Me!
5/8 x 8/15 x 3/5 x 1/2
18&3/8 - 12&1/24

Answer by AnlytcPhil(1116)

5/8 x 8/15 x 3/5 x 1/2 

 5     8     3     1
——— × ——— × ——— × ———
 8    15     5     2

Cancel the 8's

       1
 5     8     3     1
——— × ——— × ——— × ———
 8    15     5     2
 1


Cancel the 5's

 1     1
 5     8     3     1
——— × ——— × ——— × ———
 8    15     5     2
 1           1

Cancel the 3 into the 15 and get 5

 1     1     1
 5     8     3     1
——— × ——— × ——— × ———
 8    15     5     2
 1    5      1


Multiply all numerators and denominators:

             1
           ————
            10

----------------------------------------
----------------------------------------

   18 3/8 - 12 1/24 


  18 3/8   =    18 9/24 
- 12 1/24  =  - 12 1/24
—————————     —————————
                 6 8/24 = 6 1/3


Edwin McCravy
AnlytcPhil@aol.com